What amount of money invested today at 3.54% compounded semiannually will have an accumulated value of $551,000 in 7 years from now. Round all answers to two decimal places if necessary.


1
Expert's answer
2022-01-24T15:59:18-0500

Solution:

"A=\$551,000; t=7;n=2;r=3.54\%"

Let the invested money be $P.

Using "A=P(1+\dfrac{r}{100n})^{nt}"

"\Rightarrow 551000=P(1+\dfrac{3.54}{200})^{14}
\\ \Rightarrow 551000=P(1.2784)
\\ \Rightarrow P=\$431007.51"